# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v8

@@ 6 Using the Data Structure
Having defined the GLOVER data structure, we are ready to begin using it. The first and most important concept that we will define that will guide our research is the decreasing sequence property. With that and the GLOVER data structure, and the ammunition of knowing that any oriented graph without back-arcs can be represented in a GLOVER, we can assume we are working in this environment.
The decreasing sequence property explains the nature of how the exterior degrees, and thus the sizes of the containers, decrease as we get further from $v_{0}$. Exterior degrees, though, are a part of a node’s degree equation. There’s another part of it, though. As we move further from $v_{0}$, if there are no back arcs, the size of the interior degrees must increase.
This phenomenon reflects a yin and yang relationship between the interior and exterior degrees. In all of this, the minimum degree node, $v_{0}$, is serving as a catalyst and driving force for the process. At the other end of the spectrum, the decreasing sequence property is the bottleneck limiting the expansion of the graph’s out-degrees.
The case where $d^{+}(v_{0})=3$ and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) are important because they provide information about what a counterexample to the conjecture would involve. A graph $G$ was constructed in which the out-degree of $v_{0}$ does not double in the square $G^{2}$. A simple graph of degree three corresponds to the smallest cycle that can exist in an undirected graph. Therefore, if the edges are oriented in the same direction, a cycle of order three can be created among the container adjacent to $v_{0}$. If we look at this situation further, we see that in this example, a node that is an out-neighbor of $v_{0}$ had its out-degree (at least) double. This section will initiate the examination of the adequate conditions for other nodes in the graph to not have their degrees double.
@@
Two competing new ideas have been presented: the decreasing sequence property (DSP) and the minimal degree node. A crucial starting point for our approach and degree values is a minimal degree node. It acts as the anchor by being chosen as the beginning point. The remaining nodes are currently being forced into the attributes that will surface in the counterexample construction by the minimal degree node. It acts as a catalyst, influencing other nodes and setting up the conditions necessary for the subsequent analysis.
In contrast, the DSP introduces constraints on connectivity. This concept is akin to a bottleneck. A bottleneck in graph theory represents a point that limits flow or connectivity within the network[6], a role precisely filled by the DSP. Thus, the minimum degree node and the DSP together function as a catalyst and bottleneck, respectively.
@@
If we have a node $x$ with the decreasing sequence property, and $y$ is an out-neighbor of $x$, we will partition $y$’s neighbors into two sets, those that are also neighbors of $x$ and those that are not neighbors of $x$. In the proofs of Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1), we looked at the out-degree of a node inside the neighbor-induced subgraph (of $v_{0}$) and in the graph with the neighbor-induced subgraph removed. We will now give terms to these concepts.
@@
Interior Degree at least 1 Suppose we have an oriented graph $G$. Suppose further that $v_{0}$ is a minimum out-degree node for $G$ satisfying the decreasing sequence property. In an attempt to prevent $v_{0}$’s degree from doubling in $G$, every node $x$ in the neighboring container $C_{1}$ must have an interior degree of at least 1. That is, for $x\in C_{1}$,
@@
Lemma 6.2 (Interior Degree at least 1) is important because we can see some structural properties in generally oriented graphs that may coincide with degree doubling. We see that when all of the neighbors of the minimum degree node have degree 1, then it will prevent that minimum degree node from having its degree doubled.
@@
Exterior Less than Degree
Suppose we have an oriented graph $G$. If $x_{i}$ in container $C_{i}$ has the decreasing sequence property, then for all $y\in N^{+}(x_{i})|\text{ext}(x,y)|<d^{+}(x)$.
@@
Lemma 6.3 (Exterior Less than Degree) is very similar to (Exterior Less than Degree), Lemma 6.2 (Interior Degree at least 1), with the notable exception that Lemma 6.3 refers to any node with the DSP, whereas we only made Lemma 6.2 true for the minimum degree node.
@@
Andi
$\Rightarrow$ BellChadDawnElseFayeGusHugoIslaJade $\Rightarrow$ KaiLeiMiaNovaOwenPatQinRyan $\Rightarrow$ SageTateUgoVeraWadeXiaYoel $\Rightarrow$
Figure 7: This is an example of a social media setup of the SSNC. The second container has 21 arcs and seven nodes, which represents an orientation of the complete graph.
We have now defined the data structure that we will use throughout the rest of this paper. In doing so, we have a better understanding of interior and exterior arcs, which will help guide us in our search for a counterexample.
@@ 7 Back Arcs
With our discussion of the data structure concluded, we are ready to launch an investigation into back arcs. We will investigate whether these arcs can bring about a counterexample.
@@
Lemma 7.1 (Regular Interior Degrees) provides another insight into the structure within the containers of these counterexamples. It shows that all the out-neighbors of interior arcs within a container must have the same degree. This may not be surprising, as the examples we have been showing have all had this property. For example, both containers 1 and 2 in example 6.3 are orientations of regular graphs. The second container is actually an orientation of $K_{5}$, the complete graph on 5 nodes. This result begins to lend itself to some of the mathematical symmetry and structure that must be present in these containers and the SSNC.
@@
The assumptions for this letter are that we have one parent node whose interior degree doubles, and it maps the two child nodes that have different interior degrees. What happens is that because these child nodes have different interior degrees, it causes this pair node to have an exterior degree that doubles by the week exterior decreasing lemma.
@@
This means that the back arc will double $u_{i}$’s exterior degree. Then, because we’ve already assumed that $u_{i}$’s interior degree will double, we can say that $u_{i}$’s overall degree will double. This is true because its interior degree and its exterior degree will double. ∎
The proof above demonstrates that a node cannot have a back arc to an early container. To understand this proof, we must first investigate the assumptions. Here, we are assuming two things. That the node $u_{i}$ has its interior degree doubled and that the first back arc from $u_{i}$ is to the node $v_{k}$. As we saw in the proof, $v_{k}$ has more exterior neighbors than you survive, which is what causes $u_{i}$’s degree to double. The question is whether we need the condition of $u_{i}$ interior degree doubling. The weak decreasing exterior lemma tells us that we have a relationship between the exterior degrees of use by and $v_{k}$, not the overall degrees of these nodes. By assigning the interior degree to double, we can leverage this lemma.
@@
Back Arcs Special Case
  1. 1.
    Suppose $x\in C_{i}$ and $y\in C_{i+1}$, both satisfying the Decreasing Sequence Property (DSP).
  2. 2.
    The interior degree of $y$ with respect to $x$ doubles in $G[N^{+}(x)]^{2}$.
  3. 3.
    There exists a $w\in\text{ext}(x,y)$ such that the $int(y,w)$ doubles.
Then, for all $z\in\text{ext}(x,y)$, we have:
$|\text{ext}(y,z)|<|\text{ext}(x,y)|$
First, we must notice that the order within the container system, which defines exterior arcs as going to the next container (i.e., from container $C_{i}$ to container $C_{i+1}$) and not two containers down (i.e., not from container $C_{i}$ to container $C_{i+2}$), explicitly forbids $x$ having an exterior neighbor $z$ in the container $C_{i+2}$ where $x$ is in container $C_{i}$. This gives the only possibility for $z$ as being an interior neighbor of $x$.
Consequently, the arc $w,z$ is an early back arc as we saw in 7.2 (No Back Arc to Early Container). These arcs can lead to double counting. An out-neighbor of $w$ might be an out-neighbor of $x$ and a second out-neighbor of $y$. This implies that under the assumption that $w$’s interior degree doubles, $w$’s overall degree will double through that early back arc.
Lemma 7.4.
First Back Arc Between Consecutive Containers in an Oriented Graph: Let $G$ be an oriented graph with minimum degree node $v_{0}$, with minimum degree $d$. Suppose that all arcs point forward (i.e., from container $C_{j}$ to container $C_{j+1}$ for $j<i$) with no back arcs between containers $C_{1},C_{2},\dots,C_{i-1}$. If the first back arc occurs from container $C_{i}$ to $C_{i-1}$, then it will introduce an arc violating $G$’s oriented property.
@@
Lemma 7.5.
No Back Arcs Non-Existence of Back Arc In an oriented graph $G$, with a minimum degree node $v_{0}$. Let containers $C_{1}$, $C_{2}$, $\dots$, and $C_{k}$. Then there cannot exist back arcs between a node $u_{i}\in C_{i}$ and a node $v_{j}\in C_{j}$, where $j<i$.
Proof.
This was proved in the lemmas 7.2 (No Back Arc to Early Container) and 7.4 (First Back Arc Between Consecutive Containers). ∎
@@ 8 Interior Degrees Doubling
This section introduces the concept of interior degree-doubling. This is like solving independent subproblems, where the subproblems we are attempting to solve here are the doubling of a node’s interior degree inside of a container. This can be seen as a localized version of Seymour’s Second Neighborhood Conjecture, but for individual containers instead of the overall graph.
Example 8.1.
[…]
Figure 9: Illustration of the proof of 8.1 (Interior Degrees Double). This is another illustration of the Interior Degree Doubling. Each node inside a container represents a node whose degree is doubled by the inductive hypothesis. This process cannot continue forever, so we must reach the last node, and the parent of this last node in the sequence is the node whose exterior degree will double.
@@
Interior Degrees Double Let $G$ be an oriented graph. Let $v_{0}$ be a minimum-degree node in $G$ where all nodes satisfy the decreasing sequence property (DSP). Then, at least one node in each container $C_{i}$ will have a node with its interior degree doubled.
@@
Lemma 8.1 (Interior Degrees Double) is a local version of Seymour’s Second Neighborhood Conjecture. We were able to isolate the containers as an oriented graph and searched for a node whose (interior) degree doubles. We were able to set up a GLOVER data structure inside this local version starting with a minimum degree node of the container. We can then form local containers that are each their own respective distance from the minimum degree node. By strong induction, we have that each of these containers will have a node whose interior degree doubles. Then by the weak decreasing exteriors lemma, we have a decreasing set of exteriors inside this container. When this sequence gets down to size two, that node will have all exterior neighbors. Therefore, a parent of that node will have its degree doubled.
@@
The way the proof was iterated was through both strong induction and a monotonically decreasing sequence of subsets. This cannot continue indefinitely. This gives rise to why the Decreasing Sequence Property (DSP) received its name. Prior to proofs, our research showed the Weak Decreasing Exteriors Lemma as an early observation through many examples. The question is whether this was always true or whether back arcs like lemma 7.3 could force this into being a false positive.
This proof also helps with explaining the concept of the SSNC. Any oriented graph in which we want to seek a degree-doubling node can be represented in a GLOVER, as long as there are no back arcs. Once we have done this, there are containers spread out from a chosen minimum degree node. Now for each container, we can simply run the conjecture on that container. The difference is that with lemma 8.1 (Interior Degrees Double), we already know that there is a node whose degree doubles inside that container. Then with lemma 9.2 (Indistinguishable Regular Interior Nodes), we know that any node we select in that container will also have its interior degree doubled.
@@ 9 Container Density
Now that we have established interior regularity and interior degree doubling, we can start to think about what those things mean. The Weak Decreasing Exteriors Lemma implies that containers will decrease in size. We need to see how that decrease will impact the degrees of the nodes, the other properties of containers, and the development of our algorithm.
@@
Because Lemmas 8.1 (Interior Degrees Doubling) and 7.5 (No Back Arcs) both hold, we can apply Lemma 6.4 (Weak Decreasing Exteriors Lemma) here. This says that with each increase in distance from the minimum degree node, our exterior size is decreased by 1.
@@
This proof begins with two nodes in the same container. By that assumption and the regularity lemma on interior degrees, we know that these nodes have the same interior degree. Our next assumption is that these two nodes have different exterior degrees. We simply want to observe what effect that has on a parent of these nodes. The irregular exterior degree causes a ripple in the decrease in exteriors, which should be guaranteed by the weak Decreasing Exteriors Lemma. The container $C_{i+1}$ is now the same size as $C_{i}$. This means that a node in $C_{i-1}$ will have their exterior degree double.
This is an important property because it ties into the Decreasing Sequence Property (DSP). We were able to use the 6.4 (Weak Decreasing Exteriors Lemma) in this proof because the two go hand in hand.
Anything that leads to a hiccup in the decrease in exteriors would cause a node’s degree to double. The DSP is a bottleneck in this way because it is really limiting what can happen in this graph in the instance of not having a node’s degree double.
@@
In Section 5 (Building a Data Structure), we demonstrated that mathematical induction could be applied to our framework. This proof highlights the utility of the lemma. It allowed us to construct a proof based on the container distance from the chosen minimum-degree node. It is possible that this result could have been proven without relying on induction. We could have used a different approach or simply avoided induction based on container distance from $v_{0}$ entirely. Still, the use of induction provides an important method, particularly for graphs, where the usage of induction is more limited. Induction may not always be necessary, but having a variety of alternative approaches to proofs strengthens our theoretical foundation and enhances our flexibility in tackling these problems.
Similar to the proof of lemma 6.4 (Weak Decreasing Exteriors Lemma), we assume here that there is a node in each container whose degree doubles. The difference is that that assumption does not need to be explicitly stated because we have already proved that there is a node whose interior degree doubles in each container. Similarly, this lemma requires there to be no back arcs, like the weak decreasing exteriors lemma. We’ve already proved that back arcs cannot exist without creating a degree doubling node, though. As a result, we do not need to state that assumption.
@@
Example 9.1 shows some of the patterns that emerge with this regularity in the SSNC. Starting from the outside in, going inward represents the order of containers as they increase in distance from $v_{0}$. This example has visualized only the interior arcs, as the exterior arcs would complicate this visualization. However, just as in previous examples, all nodes in a previous container are mapping to all nodes in the next container.

### v9

@@ 6 Using the Data Structure
6.1 Decreasing Sequence Property
Having laid the ground work for the GLOVER data structure, we are ready to begin using it. The first and most important concept that we will define that will guide our research is the decreasing sequence property. With that and the GLOVER data structure, and the ammunition of knowing that any oriented graph without back-arcs can be represented in a GLOVER, we can assume we are working in this environment.
The decreasing sequence property explains the nature of how the exterior degrees, and thus the sizes of the containers, decrease as we get further from $v_{0}$. Exterior degrees, though, are a part of a node’s degree equation. Still, there’s another part of it. As we move further from $v_{0}$, if there are no back arcs, the size of the interior degrees must increase.
This phenomenon reflects a yin and yang relationship between the interior and exterior degrees. In all of this, the minimum degree node, $v_{0}$, is serving as a catalyst and driving force for the process. At the other end of the spectrum, the decreasing sequence property is the bottleneck limiting the expansion of the graph’s out-degrees. The case where $d^{+}(v_{0})=3$ and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) are important because they provide information about what a counterexample to the conjecture would involve. A graph $G$ was constructed in which the out-degree of $v_{0}$ does not double in the square $G^{2}$. A simple graph of degree three corresponds to the smallest cycle that can exist in an undirected graph. Therefore, if the edges are oriented in the same direction, a cycle of order three can be created among the container adjacent to $v_{0}$. If we look at this situation further, we see that in this example, a node that is an out-neighbor of $v_{0}$ had its out-degree (at least) double. This section will initiate the examination of the adequate conditions for other nodes in the graph to not have their degrees double.
@@ 6.1 Decreasing Sequence Property
Two competing new ideas have been presented: the decreasing sequence property (DSP) and the minimal degree node. A crucial starting point for our approach and degree values is the minimal degree node $v_{0}$. It acts as the anchor by being chosen as the starting point. The remaining nodes are currently being forced into the attributes that will surface in the counterexample construction by $v_{0}$. It acts as a driving force, influencing other nodes and setting up the conditions necessary for the subsequent analysis.
In contrast, the DSP introduces constraints on connectivity. This concept is akin to a bottleneck. A bottleneck in graph theory represents a point that limits flow or connectivity within the network [5], a role precisely filled by the DSP. Thus, the minimum degree node $v_{0}$ and the DSP property together function as a catalyst and bottleneck, respectively.
@@
If we have a node $x$ with the decreasing sequence property, and $y$ is an out-neighbor of $x$, we will partition $y$’s neighbors into two sets, those that are also neighbors of $x$ and those that are not neighbors of $x$. In the proofs of Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1), we looked at the out-degree of a node inside the neighbor-induced subgraph (of $v_{0}$) and in the graph with the neighbor-induced subgraph removed. The neighbors of $y$ that are also neighbors of $x$ are called interior neighbors, while the neighbors of $y$ that are not neighbors of $x$ are called exterior neighbors. One thing we see in Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) is that the interior neighbors of $v_{0}$ helped $v_{0}$’s degree not to double. They did this by all having a degree of 1.
@@
Interior Degree at least 1 Suppose we have an oriented graph $G$. Suppose further that $v_{0}$ is a minimum out-degree node for $G$. In an attempt to prevent $v_{0}$’s degree from doubling in $G$, every node $x$ in the neighboring container $C_{1}$ must have an interior degree of at least 1. That is, for $x\in C_{1}$,
@@
Lemma 6.2 (Interior Degree at least 1) is important because we can see some structural properties in generally oriented graphs that may coincide with degree-doubling. We see that when all of the neighbors of the minimum degree node have degree 1, then it will prevent that minimum degree node from having its degree-doubled. This is a generalization of Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1).
@@
Exterior Less than Degree Suppose we have an oriented graph $G$. If $x_{i}$ in container $C_{i}$ has the decreasing sequence property, then for all $y\in N^{+}(x_{i})|\text{ext}(x,y)|<d^{+}(x)$.
@@
Lemma 6.3 (Exterior Less than Degree) is very similar to lemma 6.2 (Interior Degree at least 1), with the notable exception that Lemma 6.3 refers to any node with the DSP, whereas we only made Lemma 6.2 true for the minimum degree node.
@@ 6.2 Representation
Figure 7: This is an example of a social media setup of the SSNC. The third container has 21 arcs and seven nodes, which represents an orientation of the complete graph.
We have now defined the data structure that we will use throughout the rest of this paper. In doing so, we have a better understanding of interior and exterior arcs, which will help guide us in our search for a counterexample. We also have an understanding of the Decreasing Sequence Property which will guide the nodes throughout the rest of this proof process.
@@ 7 Back Arcs
With our discussion of the data structure concluded, we are ready to launch an investigation into back arcs. We will investigate the influence these back arcs have on our data structure, and our search for a counterexample.
@@
Lemma 7.1 (Regular Interior Degrees) provides another insight into the structure within the containers of these counterexamples. It shows that all the out-neighbors of interior arcs within a container must have the same degree. This may not be surprising, as the examples we have been showing have all had this property. For example, both containers 1 and 2 in example 6.3 are orientations of regular graphs. The second container is actually an orientation of $K_{7}$, the complete graph on 7 nodes. This result begins to lend itself to some of the mathematical symmetry and structure that must be present in these containers and the SSNC.
@@
The assumptions for this lemma are that we have one parent node whose interior degree doubles, and it maps the two child nodes that have different interior degrees. What happens is that because these child nodes have different interior degrees, it causes this parent node to have an exterior degree that doubles by lemma 6.4 (Weak Decreasing Exteriors Lemma).
@@
$v_{k}$’s neighbors are distinct from $u_{i}$’s because it is in a different container. We formed these containers by a partition, so these containers cannot overlap. This means that the back arc will double $u_{i}$’s exterior degree. Then, because we’ve already assumed that $u_{i}$’s interior degree will double, we can say that $u_{i}$’s overall degree will double. This is true because its interior degree and its exterior degree will double. ∎
The proof above demonstrates that a node cannot have a back arc to an early container. To understand this proof, we must first investigate the assumptions. Here, we are assuming two things. That the node $u_{i}$ has its interior degree doubled and that the first back arc from $u_{i}$ is to the node $v_{k}$. As we saw in the proof, $v_{k}$ has more exterior neighbors than $u_{i}$, which is what causes $u_{i}$’s degree to double. The question is whether we need the condition of $u_{i}$ interior degree doubling. Lemma 6.4 (Weak Decreasing Exteriors Lemma) tells us that we have a relationship between the exterior degrees of $u_{i}$ and $v_{k}$, not the overall degrees of these nodes. By assigning the interior degree to double, we can leverage this lemma.
@@
First Back Arc Between Consecutive Containers in an Oriented Graph: Let $G$ be an oriented graph with minimum degree node $v_{0}$, with minimum degree $\delta$. Suppose that all arcs point forward (i.e., from container $C_{j}$ to container $C_{j+1}$ for $j<i$) with no back arcs between containers $C_{1},C_{2},\dots,C_{i-1}$. If the first back arc occurs from container $C_{i}$ to $C_{i-1}$, then the first back arc will introduce an arc violating $G$’s oriented property.
@@
Example 7.2.
Back Arcs Special Case
  • Suppose $x\in C_{i}$ and $y\in C_{i+1}$.
  • $w$ is a second out-neighbor of $x$ and a first out-neighbor of $y$.
  • $z$ is an out-neighbor of $w$, and an out-neighbor of $x$ and an out-neighbor of $y$.
  • This implies that $z$ is in container $C_{i+1}$, the same container as $y$.
  • By 7.1 (Regular Interior Degrees), we know that because the nodes in container $C_{i+1}$ cannot have back arcs, they can only have interior and exterior arcs.
  • This means that by 7.3 (No Back Arc Between Consecutive Containers) that every node (including $y$ and $z$ in $C_{i+1}$ will map to the same set of exterior nodes in container $C_{i+2}$.
  • Thus, $z$ cannot be a back arc of $w$ and an exterior arc of $y$.
Example 7.3.
[…]
The concern with 6.4 (Weak Decreasing Exteriors Lemma) concern was about the potential for overlapping neighborhoods when considering how interior and exterior degrees change. If a node $w$ is both an interior neighbor (or contributes to the interior degree) and an exterior neighbor (or contributes to the exterior degree), then we couldn’t simply add the changes in interior and exterior degrees to get the change in the total degree. We’d be counting $w$’s contribution twice.
The conditions regarding back arcs and the lemma 7.1 (Regular Interior Degrees) are crucial for the validity of Lemmas 7.2 (No Back Arc to Early Container) and 7.3 (No Back Arc Between Consecutive Containers). Lemma 7.3, which is most directly relevant to the cases considered here, relies on the property of oriented graphs that if every node in container $C_{i-1}$ has an arc going to every node in container $C_{i}$, then no back arc can exist from a node in $C_{i}$ to a node in $C_{i-1}$. This property is a direct consequence of the definition of oriented graphs and the container structure. The Regular Interior Degrees Lemma (Lemma 7.1) plays a crucial role in establishing this property. It shows that if there were an imbalance in interior degrees between nodes in $C_{i}$, it would force the exterior degree of a parent node in $C_{i-1}$ to double. Combined with the Weakly Decreasing Exteriors Lemma (Lemma 6.4), which states that exterior degrees cannot increase along a path under certain conditions, this implies that interior degrees within a container must be regular (essentially the same), and thus the oriented graph property holds. The assumptions in Lemmas 7.2 and 7.3—that we are considering the first back arc—reinforce that no such back arc exists at any point in the process of constructing the containers. Therefore, the scenarios considered in these lemmas (and any example based on a back arc) are consistent with the graph structure and the regularity properties established by Lemmas 7.1 and 6.4.
@@ 8 Interior Degrees Doubling
This section introduces the concept of interior degree-doubling, a key element in our proof strategy for the SSNC. Several preceding lemmas have relied on this concept. Our overall approach is based on a divide and conquer paradigm, where each container is treated as an independent subproblem. These subproblems focus on demonstrating the doubling of a node’s interior degree within a container. This can be viewed as a localized version of the SSNC, applied to individual containers rather than the entire graph.
Figure 9: Illustration of the proof of 8.1 (Interior Degrees Double). This is a first look at a sketch of the proof of the Interior Degree Doubling. Each node inside a container represents a node whose degree is doubled by the inductive hypothesis. This process cannot continue forever, so we must reach the last node, and the parent of this last node in the sequence is the node whose exterior degree will double.
@@
Interior Degrees Double Let $G$ be an oriented graph. Let $v_{0}$ be a minimum-degree node in $G$. Then, at least one node in each container $C_{i}$ will have a node with its interior degree doubled.
@@
Lemma 8.1 (Interior Degrees Double) is a local version of Seymour’s Second Neighborhood Conjecture. We were able to isolate the containers as an oriented graph and searched for a node whose (interior) degree doubles. We were able to set up a GLOVER data structure inside this local version starting with a minimum degree node of the container. We can then form local containers that are each their own respective distance from the minimum degree node $u_{j}$. By strong induction, we have that each of these containers will have a node whose interior degree doubles. Then by lemma 6.4 (Weak Decreasing Exteriors Lemma), we have a decreasing set of exteriors inside this container. When this sequence gets down to size two, that node will have all exterior neighbors. Therefore, a parent of that node will have its degree doubled.
@@
The way the proof was iterated was through both strong induction and a monotonically decreasing sequence of subsets. This cannot continue indefinitely. This gives rise to why the Decreasing Sequence Property (DSP) received its name. Prior to proofs, our research showed lemma 6.4 (Weak Decreasing Exteriors Lemma) as an early observation through many examples. The question was whether this was always true or whether back arcs like lemma 7.2 could force this into being a false positive.
This proof also helps with explaining the concept of the SSNC. Any oriented graph in which we want to seek a degree-doubling node can be represented in a GLOVER. Once we have done this, there are containers spread out from a chosen minimum degree node. Now for each container, we can simply run the conjecture on that container. The difference is that with lemma 8.1 (Interior Degrees Double), we already know that there is a node whose degree doubles inside that container.
@@ 9 Container Density
Now that we have established interior regularity and interior degree doubling, we can start to think about what those things mean. The 6.4 (Weak Decreasing Exteriors Lemma) implies that containers will decrease in size. We need to see how that decrease will impact the degrees of the nodes, the other properties of containers, and the development of our algorithm.
@@
Because Lemmas 8.1 (Interior Degrees Doubling), 7.2 (No Back Arc to Early Container), amd 7.3 (First Back Arc Between Consecutive Containers) all hold, we can apply Lemma 6.4 (Weak Decreasing Exteriors Lemma) here. This says that with each increase in distance from the minimum degree node, our exterior size is decreased by 1.
@@
This proof begins with two nodes in the same container. By that assumption and the regularity lemma on interior degrees, we know that these nodes have the same interior degree. Our next assumption is that these two nodes have different exterior degrees. We simply want to observe what effect that has on a parent of these nodes. The irregular exterior degree causes a ripple in the decrease in exteriors, which should be guaranteed by lemma 6.4 (Weak Decreasing Exteriors Lemma). The container $C_{i+1}$ is now the same size as $C_{i}$. This means that a node in $C_{i-1}$ will have their exterior degree double.
This is an important property because it ties into the Decreasing Sequence Property (DSP). We were able to use the 6.4 (Weak Decreasing Exteriors Lemma) in this proof because the two go hand in hand. Anything that leads to a hiccup in the decrease in exteriors would cause a node’s degree to double. The DSP is a bottleneck in this way because it is really limiting what can happen in this graph in the instance of not having a node’s degree double.
@@
In Section 5 (Building a Data Structure), we demonstrated that mathematical induction could be applied to our framework. This proof highlights the utility of the lemma. It allowed us to construct a proof based on the container distance from the minimum-degree node $v_{0}$. It is possible that this result could have been proven without relying on induction. We could have used a different approach or simply avoided induction based on container distance from $v_{0}$ entirely. Still, this use of induction provides an important method, particularly for graphs, where the usage of induction is more limited. Induction may not always be necessary, but having a variety of alternative approaches to proofs strengthens our theoretical foundation and enhances our flexibility in tackling these problems.
Similar to the proof of lemma 6.4 (Weak Decreasing Exteriors Lemma), we assume here that there is a node in each container whose interior degree doubles. The difference is that that assumption does not need to be explicitly stated because we have already proved that there is a node whose interior degree doubles in each container. Similarly, this lemma requires there to be no back arcs, like lemma 6.4 (Weak Decreasing Exteriors Lemma). We’ve already proved that back arcs cannot exist without creating a degree-doubling node, though. As a result, we do not need to state that assumption.
@@
Example 9.1 shows some of the patterns that emerge with these interior degrees lower bound in the SSNC. Starting from the outside in, going inward represents the order of containers as they increase in distance from $v_{0}$. This example has visualized only the interior arcs, as the exterior arcs would complicate this visualization. However, just as in previous examples, all nodes in a previous container are mapping to all nodes in the next container.